The table describes hypothetical employment statistics for a country in a given year.

Labor force Full-time workers and
voluntary part-time workers
Involuntary part-time workers Discouraged workers
174 million 162 million 8 million 6 million

Based on the information, what is the unemployment rate? Enter your answer as percentage rounded to one decimal place.
Answer #1

Number of unemployed people= labor force - employed

= 174 - 162 - 8= 4 million

Unemployment rate= unemployed*100/labor force

= 4*100/174 = 2.3%